OpenStudy (anonymous):

What polynomial identity should be used to prove that 162 = (10 + 6)2? Difference of Cubes Difference of Squares Square of Binomial Sum of Cubes

OpenStudy (anonymous):

@Sepeario

OpenStudy (anonymous):

@Nnesha

rishavraj (rishavraj):

\[16^2 = (10 + 6)^2\] http://www.themathpage.com/alg/perfect-square-trinomial.htm

OpenStudy (anonymous):

thats the equation yes.

OpenStudy (sepeario):

technically he already gave you the answer..

OpenStudy (anonymous):

wouldnt it be square of binonmial?

rishavraj (rishavraj):

yup :))

OpenStudy (anonymous):

What is the simplified form of 3 over 2x minus 5 + 21 over 8 x squared minus 14x minus 15 ? 6 times the quantity 2 x plus 5 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 times the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 time the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 times the quantity x plus 1 end quantity over the quantity 2x plus 5 end quantity times 4 x plus 3

OpenStudy (anonymous):

wht about this

rishavraj (rishavraj):

\[\frac{ 3 }{ 2x - 5 } + \frac{ 21 }{ 8x^2 - 14x - 15 }\] u see 8x^2 - 14x - 15 = (2x - 5)(4x + 3)

OpenStudy (anonymous):

yes. now what

OpenStudy (anonymous):

hello?